OpenStudy (anonymous):

An inequality is shown below: -np - 4 ≤ 2(c - 3) Which of the following solves for n? n ≤ - the quantity 2 times c minus 2 all over p n ≤ - the quantity 2 times c minus 10 all over p n ≥ - the quantity 2 times c minus 2 all over p n ≥ - the quantity 2 times c minus 10 all over p

OpenStudy (anonymous):

hello ranga can you help me with this question

OpenStudy (ranga):

-np - 4 ≤ 2(c - 3) Wee need to solve for n. That means isolate n. add 4 to both sides: -np ≤ 2(c - 3) + 4 -np ≤ 2c - 6 + 4 -np ≤ 2c - 2 divide by p -n ≤ (2c - 2) / p divide by -1. And because it is a negative number, flip the inequality n >= -(2c - 2) / p

OpenStudy (anonymous):

thankyou do you think you could help me with some other questions

OpenStudy (ranga):

Post one more here and i can try.

OpenStudy (anonymous):

okay hold on

OpenStudy (anonymous):

A hybrid car can travel 40 miles per gallon. Approximately how many gallons of fuel will the car need to travel 20 km? [1 mile = 1.6 km] 0.3 0.8 2.5 3.2

OpenStudy (ranga):

Convert 20 km to miles. (You can also convert 40 miles to km. Either way you will get the same answer). 1 miles = 1.6 km. How many miles is 20 km. Proportions problem. x miles / 20 km = 1 mile / 1.6 km x = 20 * 1 / 1.6 = 12.5 miles. 40 miles per gallon = 40 miles / 1 gallon. How many gallons will 12.5 miles need? Once gain, proportions problem. 12.5 miles / x gallons = 40 miles / 1 gallon x = 12.5 * 1 / 40 = ? gallons

OpenStudy (anonymous):

ok thankyou could you help with just one more question

OpenStudy (ranga):

okay.

OpenStudy (anonymous):

thanks so much

OpenStudy (anonymous):

Martin combined different amounts of orange juice, pineapple juice, and strawberry juice to make 1850 milliliters (ml) of fruit punch. The amount of orange juice combined was x ml. The amount of pineapple juice combined was 100 ml more than the amount of orange juice, and the amount of strawberry juice was half the amount of orange juice. This is represented in the equation below. x + (x + 100) + one half x = 1850 What was the difference in the amounts of pineapple juice and strawberry juice Martin combined? 100 ml 450 ml 700 ml 950 ml

OpenStudy (ranga):

x + (x + 100) + x/2 = 1850 Solve for x first. Then, Pineapple juice: x + 100 Strawberry juice: x/2 Difference: Subtract x + 100 - x/2 = x/2 + 100. Put the value of x into x/2 + 100.

OpenStudy (anonymous):

thank you so much

OpenStudy (anonymous):

for your time merry christmas

OpenStudy (ranga):

Merry Christmas and Happy Holidays to you as well. You are welcome.